OpenStudy (anonymous):

Two triangles are similar and have a ratio of similarity of 1:5. What is the ratio of their perimeters and the ratio of their areas? perimeter: 1:5 area: 1:25 perimeter: 1:5 area: 1:10 perimeter: 10:1 area: 5:1 perimeter: 25:1 area: 5:1

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

perimeter will have the same ratio area will have the square of the original ratio

OpenStudy (anonymous):

go with the first one

OpenStudy (anonymous):

1 to 5 and 1 to 5^2 which is: 1 to 5 and 1 to 25 first answer

OpenStudy (anonymous):

This comes from: Perimeter of the first triangle = x + y + z Perimeter of the second triangle = 5x + 5y + 5z = 5(x + y + z) Ratio = (x + y + z) / 5(x + y + z) = 1/5

OpenStudy (anonymous):

uw! @InsanelyChaotic

OpenStudy (anonymous):

The second ratio of 1/25 comes from: 1st triangle: area = (1/2)(b)(h) 2nd triangle: area = (1/2)(5b)(5h) = 25[(1/2)(b)(h)] Ratio = (1/2)(b)(h) / 25[(1/2)(b)(h)] = 1/25